OpenStudy (anonymous):

{(x,y) : x-3y=4} x-3y=4 2x-6y=Q Solve for Q? Are you supposed to use elimination? Is that possible without Q?

OpenStudy (anonymous):

if we use elimination, we should make x or y the same size before subtracting

OpenStudy (anonymous):

When I do that, -2(-2x+6y=-8) so -2x+6y=-8 2x-6y=Q Both variables cross out, right?

OpenStudy (anonymous):

exactly; so we get something like 0=-8+Q

OpenStudy (anonymous):

How?

OpenStudy (anonymous):

add the two lines: -2x+6y=-8 2x-6y=Q ------------- -2x+2x +6y-6y = -8+Q 0x + 0 y = -8+Q